Graph the equation using the x and y intercepts.
x/3-y/7=1

Answer by Alan3354(69443)

Graph the equation using the x and y intercepts.
x/3-y/7=1
==========
A = x-int = (3,0)
B = y-int = (0,-7)
Plot the points, draw a line thru them.
------
Also,
Bx + Ay = A*B
-7x + 3y = -21
7x - 3y = 21